OpenStudy (heart_offire):

need help x + 3y = 7 x - 3y = 1 Solve the system of equations. A) x = 4, y = 1 B) x = 1, y = 4 C) x = -2/3, y = 3 D) x = 3, y = -2/3

OpenStudy (heart_offire):

WAS

OpenStudy (mathmale):

Take a good look at the 2 equations given. I see a very easy way in which youi could solve this system. But first, what would YOU do?

OpenStudy (heart_offire):

maybe cancellation

OpenStudy (mathmale):

Yes. Add the 2 equations together. Please share your result.

OpenStudy (heart_offire):

after i cancel all i get is x=7 and x=1

OpenStudy (mathmale):

I'd like to see your results, please (including all steps you took to combine these equations).

OpenStudy (mathmale):

x + 3y = 7 x - 3y = 1 You have to combine all terms, column by column. for example, combining the 2nd column results in 0 for that column. Please combine columns 1 and 3 now.

OpenStudy (heart_offire):

so 7 +1 =8

OpenStudy (heart_offire):

and that would be column 3

OpenStudy (heart_offire):

and column 1would just be x

rebeccaxhawaii (rebeccaxhawaii):

it would be 2x=8

OpenStudy (mathmale):

"and column 1would just be x" Not so. You end up with 2x=8. Solve this for x. Next, subst. your newly-found x-value into either of the given equations to find y. Write your solution in the form (x, y).

OpenStudy (heart_offire):

4,1